# EPO: Entropy-regularized Policy Optimization for LLM Agents Reinforcement Learning

### v1

@@ Abstract
Training LLM agents in multi-turn environments with sparse rewards, where completing a single task requires 30+ turns of interaction within an episode, presents a fundamental challenge for reinforcement learning. We identify a critical failure mode unique to this setting: the exploration-exploitation cascade failure. This cascade begins with early-stage policy premature convergence, where sparse feedback causes agents to commit to flawed, low-entropy strategies. Subsequently, agents enter late-stage policy collapse, where conventional entropy regularization becomes counterproductive, promoting chaotic exploration that destabilizes training. We propose Entropy-regularized Policy Optimization (EPO), a general framework that breaks this failure cycle through three synergistic mechanisms: (1) adopting entropy regularization in multi-turn settings to enhance exploration, (2) an entropy smoothing regularizer that bounds policy entropy within historical averages to prevent abrupt fluctuations, and (3) adaptive phase-based weighting that balances exploration and exploitation across training. Our analysis justifies that EPO guarantees monotonically decreasing entropy variance while maintaining convergence. EPO achieves up to 152% performance improvement on ScienceWorld and up to 19.8% on ALFWorld. Our work demonstrates that multi-turn sparse-reward settings require fundamentally different entropy control than traditional RL, with broad implications for LLM agent training. The code is available at the URL  [^1] https://github.com/WujiangXu/EPO.
@@ 1 Introduction
Reinforcement Learning (RL) (Shao et al., 2024; Schulman et al., 2017) has become an important approach for post-training Large Language Models (LLMs), enabling LLMs to acquire reasoning ability with single-turn verified reward (Guo et al., 2025; Yang et al., 2025). LLM agents start to achieve success in versatile scenarios, such as coding (Anthropic, 2025; Kipruto & Salva, 2025; OpenAI, 2025b; Wang et al., 2025b), computing using (OpenAI, 2025a; Xie et al., 2024; Zheng et al., 2024), and searching (Team et al., 2025; Zeng et al., 2025), benefiting from LLM’s powerful reasoning ability. Multi-turn LLM agent scenarios present unique challenges compared to traditional LLM RL settings, as episodes can span 30+ turns (Shridhar et al., 2021; Wang et al., 2022) with extended trajectories and sparse rewards. Unlike conventional RL, where rewards are distributed throughout the episode, LLM agents typically receive feedback only at completion, meaning the policy is updated based on the shared reward across long-horizon turns within an episode, making the exploration-exploitation trade-off (Sutton, 1988) crucial for stable training.
Figure 1: The exploration-exploitation cascade failure in multi-turn agent training. The cascade failure manifests in two distinct phases clearly visible in the figure: (1) Phase 1 - Excessive Early Exploration (0-40 steps): PPO’s early trajectory steps (pink dashed line) exhibit rapid, uncontrolled entropy growth that creates unstable behavioral foundations, while rewards remain stagnant, indicating ineffective exploration-to-reward conversion. (2) Phase 2 - Uncertainty Propagation (40-120 steps): The instability from early steps cascades to late trajectory steps (red dotted line), maintaining dangerously high entropy oscillations that prevent coherent strategy formation and result in reward plateaus despite continuous exploration. This two-phase pattern demonstrates why standard entropy methods fail in multi-turn environments. In contrast, our EPO method maintains stable, controlled entropy levels across both early and late trajectory steps throughout training, achieving significantly lower final entropy values and consistent reward improvement, preventing the cascade failure.
Traditional RL approaches (Haarnoja et al., 2018; Mnih et al., 2016; Williams & Peng, 1991) employ entropy regularization by adding the entropy of the policy to the objective function, discouraging premature convergence to suboptimal deterministic policies through a hyperparameter-controlled regularization term. In RL training for LLM, researchers have adapted these mechanisms by incorporating entropy into the advantage function to reward high-entropy token generation (Cui et al., 2025; Dong et al., 2025; He et al., 2025; Wang et al., 2025a), encouraging exploration and preventing entropy collapse in later training stages. However, entropy-controlled LLM methods prove fundamentally inadequate for multi-turn agent environments because they cannot resolve the exploration-exploitation cascade failure. This failure cascade proceeds in two distinct phases, as illustrated in Figure 1. First, sparse and delayed reward signals induce excessive early-stage exploration, where standard entropy regularization causes uncontrolled entropy growth in early trajectory steps rather than premature convergence to low-entropy states. This blind exploration creates unstable foundations that systematically commit to suboptimal behavioral patterns. Having locked into these flawed early trajectories, the agent then enters late-stage uncertainty propagation, where accumulated uncertainty from early steps compounds in late steps, maintaining dangerously high entropy levels that prevent coherent strategy formation and further degrade performance (seen from PPO’s persistently high entropy in  Figure 1). In contrast, our EPO method controls cross-step entropy propagation, maintaining stable entropy levels and achieving consistent reward improvement. Therefore, how to design exploration mechanisms that navigate the exploration-exploitation tradeoff without triggering cascade failure remains an important problem for multi-turn agent training.
@@
To address this cascade issue, we propose Entropy-regularized Policy Optimization (EPO), a novel framework that combines entropy regularization with specialized mechanisms for stable on-policy training under sparse reward conditions. Our framework is guided by the central insight that standard entropy regularization is insufficient because it lacks temporal awareness. We find that anchoring the policy’s entropy to a dynamically adjusted historical bound provides the necessary stability to halt the exploration-exploitation cascade failure without sacrificing essential exploration.
Our contributions are fourfold. ❶ We identify and formally characterize the exploration-exploitation cascade failure in multi-turn sparse-reward environments, revealing how excessive early-stage exploration leads to unstable behavioral foundations that compound into late-stage uncertainty propagation, a phenomenon that standard entropy methods cannot address. ❷ We adapt entropy regularization to the multi-turn setting by computing entropy across all turns within trajectories and averaging over trajectory batches, providing a foundation that captures the unique temporal structure of agent interactions. ❸ We introduce an entropy smoothing regularizer that penalizes deviations from historical entropy averages, effectively dampening the severe oscillations between overconfidence and over-exploration that plague sparse reward environments. ❹ We develop an adaptive weighting scheme that dynamically balances exploration and exploitation across training phases, starting with conservative exploration, transitioning through balanced exploration-exploitation, and ending with strong stabilization to ensure convergence. Together, these components create a theoretically grounded and general framework that guarantees monotonically decreasing entropy variance and optimal exploration-exploitation trade-offs while being compatible with any on-policy optimization method. We validate EPO on challenging benchmarks ScienceWorld (Wang et al., 2022) and ALFWorld (Shridhar et al., 2021), achieving up to 152% performance improvement with significantly more stable training dynamics, transforming previously untrainable sparse-reward scenarios into smoothly converging optimization problems.
@@ 2.1 Reinforcement Learning for LLMs
RLHF (Ouyang et al., 2022) and DPO (Rafailov et al., 2023) have become foundational approaches for aligning LLMs with human preferences, with both methods significantly improving model alignment and instruction-following capabilities. Recent RL methods such as GRPO (Shao et al., 2024) and DAPO (Yu et al., 2025) further enhance LLM reasoning abilities during post-training through verified rewards. In contrast to PPO (Schulman et al., 2017), these methods leverage batch-wise advantage computation from identical prompts, obviating the critic model and substantially improving the computational tractability of large-scale RL training for LLMs. However, a primary challenge in applying RL to LLMs is the phenomenon of policy entropy collapse: models rapidly reduce their stochasticity, converging on narrow, over-optimized behaviors (Cui et al., 2025; Dong et al., 2025; Wang et al., 2025a; Deng et al., 2025; He et al., 2025; Cheng et al., 2025a; b). In response, recent work has focused on integrating entropy control mechanisms into the optimization process to preserve policy diversity. For instance, Cui et al. (Cui et al., 2025) regulate the impact of high-covariance tokens by applying a clipping function and a KL penalty. Meanwhile, Cheng et al. (Cheng et al., 2025b) augment the standard advantage function with a clipped, gradient-detached entropy term, which encourages deeper reasoning chains without altering the original policy optimization direction.
@@ 2.2 Reinforcement Learning for LLM Agents
To enhance their autonomy, LLM agents are designed to interact with external environments using diverse toolsets (OpenAI, 2025a; Team, 2025). However, training agents to complete multi-step tasks with these tools presents significant challenges for standard reinforcement learning, including sparse rewards and credit assignment problems. To address these issues, seminal works have introduced advanced training paradigms such as hierarchical RL (Zhou et al., 2024), autonomous learning (Bai et al., 2024), and off-policy Q-learning (Bai et al., 2025). Meanwhile, another line of research employs supervised fine-tuning (SFT) to directly enhance the models’ decision-making abilities, training them on vast datasets of high-quality tool-use trajectories to master complex environments and APIs (Xi et al., 2024; Zhang et al., 2024; Qin et al., 2024). Recently, to leverage the training stability of GRPO (Shao et al., 2024), researchers have extended single-turn GRPO to multi-turn training settings with various training techniques to improve performance (Jin et al., 2025; Feng et al., 2025; Wang et al., 2025c). To guide learning in long-horizon scenarios, RLVMR (Zhang et al., 2025) introduces verifiable meta-reasoning rewards that provide dense, intermediate feedback on the agent’s reasoning process. However, these methods overlook the critical exploration-exploitation dynamics unique to multi-turn settings, failing to address the exploration-exploitation cascade failure where policies undergo early-stage premature convergence followed by late-stage excessive exploration with high entropy that makes training vulnerable to noise. This fundamental gap leads to unstable training dynamics and prevents agents from effectively learning long-horizon strategies.
@@ 3.1 On-policy Optimization
where $A^{\pi_{\theta}}(s_{t},a_{t})=Q^{\pi_{\theta}}(s_{t},a_{t})-V^{\pi_{\theta}}(s_{t})$ is the advantage function. While the policy gradient provides an unbiased estimator of $\nabla_{\theta}J(\theta)$, directly optimizing this objective can lead to instability due to large policy updates. To address this, modern on-policy methods employ surrogate objective functions that approximate the policy gradient while ensuring stable learning. The standard policy gradient can be reformulated as the surrogate objective $L^{PG}(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta_{\text{old}}}}\left[\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}\hat{A}_{t}\right]$. However, this surrogate objective is only valid for infinitesimally small updates. Proximal Policy Optimization (PPO) (Schulman et al., 2017) addresses this limitation by constraining the policy ratio through a clipped surrogate objective:
$L^{CLIP}(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta_{\text{old}}}}\left[\min\left(r_{t}(\theta)\hat{A}_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}\right)\right]$ (2)
where $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}$ is the importance sampling ratio, $\mathbb{E}_{\tau\sim\pi_{\theta_{\text{old}}}}[\cdot]$ denotes expectation over trajectories sampled under $\pi_{\theta_{\text{old}}}$, and $\epsilon$ defines the trust region bounds. Group Relative Policy Optimization (GRPO) (Shao et al., 2024) extends PPO by modifying the advantage function computation. Instead of using standard advantages $\hat{A}_{t}$, GRPO employs group-relative advantages, computed as $\tilde{A}_{t}=\frac{R_{t}-\mu_{g}}{\sigma_{g}+\delta}$. where $R_{t}$ is the return from timestep $t$, $\mu_{g}$ and $\sigma_{g}$ are the mean and standard deviation of returns within group $g$, and $\delta$ is a small constant for numerical stability. This group-based normalization provides more stable gradient estimates.
@@ 3.2 Problem formulation
We formalize the multi-turn task as a sequential decision-making reinforcement learning problem. A single LLM agent $\pi_{\theta}$ executes a task through $T$ turns over a trajectory $\tau=(s_{0},a_{0},r_{0},\ldots,s_{T},a_{T},r_{T})$. The reward is sparse, with $r_{t}=0$ for all intermediate turns and only the final turn receiving the task outcome reward, such that the total return is $R(\tau)=\sum_{t=0}^{T}\gamma^{t}r_{t}=r_{T}$, where $r_{T}\in\{0,1\}$ represents the binary task outcome. In our experimental settings, specifically ALFWorld (Shridhar et al., 2021) and SciWorld (Wang et al., 2022), we assume an undiscounted formulation where $\gamma=1$. All turns within the same task share the final outcome reward, creating a credit assignment challenge across sequential turns. The multi-turn policy optimization differs from standard RL in that losses accumulate across all $T$ turns before parameter updates:
$L^{MT}(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta_{\text{old}}}}\left[\mathbb{E}_{t\sim T}\left[\min\left(r_{t}(\theta)A_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)A_{t}\right)\right]\right]$ (3)
where $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}$ and $A_{t}$ represents the advantage estimate at turn $t$ within the multi-turn trajectory. $\mathbb{E}_{t}[\cdot]$ denotes expectation over timesteps within turn $t$, while the outer expectation is over trajectories sampled under $\pi_{\theta_{\text{old}}}$.

### v2

@@ Abstract
Reinforcement learning has enabled LLMs to acquire reasoning through single-turn verified rewards, yet extending this paradigm to multi-turn agents—where tasks span 30+ turns with only terminal rewards—introduces fundamental challenges. We identify a critical failure mode in this setting: the exploration-exploitation cascade failure. Unlike single-turn training, multi-turn agents share policy parameters across all turns, meaning entropy adjustments cannot independently control early-turn exploration and late-turn exploitation. Combined with sparse rewards that provide no intermediate correction, this coupling causes severe entropy oscillations that destabilize training. We propose Entropy-regularized Policy Optimization (EPO), a framework that stabilizes entropy dynamics through three synergistic mechanisms: (1) trajectory-level entropy regularization that captures multi-turn temporal structure, (2) an entropy smoothing regularizer that penalizes deviations from historical averages to dampen oscillations, and (3) adaptive phase-based weighting that transitions from conservative exploration to strong stabilization across training. EPO achieves up to 152% improvement on ScienceWorld and 19.8% on ALFWorld, transforming previously unstable sparse-reward scenarios into smoothly converging optimization. Our work demonstrates that multi-turn settings require fundamentally different entropy control than traditional RL. The code is available at the URL  [^1] https://github.com/WujiangXu/EPO.
Machine Learning, ICML
@@ 1 Introduction
Reinforcement Learning (RL) (grpo; ppo) has become an important approach for post-training Large Language Models (LLMs), enabling them to acquire reasoning ability through single-turn verified rewards (dpskr1; qwen3). Extending this paradigm to multi-turn LLM agents, which interact with environments across extended episodes, has achieved promising results in coding (anthropic_claude_code; openai_codex), computer use (oaicua; osworld), and web search (kimik2; GLM4.5), yet introduces fundamental challenges. In multi-turn settings, episodes can span 30+ turns (alfworld; sciworld) yet rewards arrive only at task completion. This means early-turn actions shape entire trajectories without receiving corrective feedback, making exploration-exploitation balance (tdlearning) a critical challenge that single-turn methods handle implicitly through immediate rewards.
(a) Early Steps
(b) Late Steps
(c) PPO vs PPO+EPO
Figure 1: The exploration-exploitation cascade failure in multi-turn training. (a-b) Adjusting per-turn entropy weights fails due to shared parameters: EPO-Decay applies higher weights to early turns and lower weights to late turns, yet entropy curves remain nearly identical for both early steps (a) and late steps (b). (c) Standard PPO exhibits severe entropy fluctuation (purple) with stagnant rewards (yellow), while our proposed EPO maintains stable entropy (green, declining from 1.2 to 0.3) and achieves consistent reward improvement (red).
Traditional RL approaches (sac; a3c; 91rlentropy) employ entropy regularization to address exploration, adding the policy’s entropy to the objective function to discourage premature convergence to deterministic policies. In LLM training, researchers have adapted these mechanisms to reward high-entropy token generation (clip-conv; Rl-plus; skywork; 8020entropy), operating under the assumption that entropy collapse is the primary failure mode. Through extensive empirical analysis on standard multi-turn agent benchmarks, we uncover a critical limitation in multi-turn agent training settings: the challenge is not entropy collapse but rather uncontrolled entropy dynamics. In multi-turn training, the same policy parameters control all turns. Since all turns share parameters, adjusting entropy weights per turn cannot independently control exploration and exploitation: increasing entropy for early turns inevitably affects late turns, and vice versa. Combined with sparse rewards that provide no intermediate correction signal, this makes training dynamics inherently unstable. We term this the exploration-exploitation cascade failure.
As illustrated in Figure 1, we empirically validate this limitation. In Figure (a) and (b), we attempt a natural solution: applying higher entropy weights to early turns to promote exploration, and lower weights to late turns to encourage exploitation (EPO-Decay). However, due to shared policy parameters, early-turn and late-turn entropy curves remain nearly identical to the unweighted baseline (EPO-Base). This confirms that per-turn weight adjustment cannot break the coupling between turns. Figure (c) shows the consequence: PPO exhibits severe entropy fluctuation with stagnant rewards, while our proposed EPO maintains stable declining entropy and consistent reward growth. This cascade failure exposes a mismatch between existing entropy methods and multi-turn structure. Standard entropy regularization operates statelessly, treating each gradient update in isolation. However, in multi-turn training, shared parameters mean that each gradient update affects all turns simultaneously, so entropy at specific step depends on the entire training history, not just the current update. This raises a critical challenge: how can we stabilize entropy dynamics over time when shared parameters prevent independent per-turn control?
@@
To address this cascade issue, we propose Entropy-regularized Policy Optimization (EPO), a novel framework that combines entropy regularization with specialized mechanisms for stable on-policy training under sparse reward conditions. Our contributions are fourfold.
We empirically discover and formally characterize the exploration-exploitation cascade failure on standard multi-turn benchmarks, demonstrating that this novel failure mode, where entropy oscillations compound across turns under sparse rewards, is the primary cause of poor performance in standard RL baselines.
We extend standard entropy regularization to multi-turn settings, where entropy accumulates across sequential turns through shared parameters, requiring trajectory-level aggregation rather than per-step computation.
We introduce an entropy smoothing regularizer that penalizes deviations from historical entropy averages, effectively reducing the oscillations between overconfidence and over-exploration that characterize cascade failure.
We develop an adaptive weighting scheme that dynamically balances exploration and exploitation across training phases, applying stronger smoothing early to establish controlled exploration, relaxing mid-training as the policy stabilizes, and maintaining sufficient regularization in later phases to ensure convergence.
Together, these components create a theoretically grounded and general framework that prevents the empirically observed cascade failure and ensures optimal exploration-exploitation trade-offs while being compatible with any on-policy optimization method. We validate EPO on challenging benchmarks ScienceWorld (sciworld) and ALFWorld (alfworld), achieving up to 152% performance improvement with more stable training dynamics, transforming previously unstable sparse-reward scenarios into smoothly converging optimization problems.
@@ 2.1 Reinforcement Learning for LLMs
RLHF (rlhf) and DPO (dpo) have become foundational approaches for aligning LLMs with human preferences, with both methods significantly improving model alignment and instruction-following capabilities. Recent RL methods such as GRPO (grpo) and DAPO (dapo) further enhance LLM reasoning abilities during post-training through verified rewards. In contrast to PPO (ppo), these methods leverage batch-wise advantage computation from identical prompts, obviating the critic model and substantially improving the computational tractability of large-scale RL training for LLMs. However, a primary challenge in applying RL to LLMs is the phenomenon of policy entropy collapse: models rapidly reduce their stochasticity, converging on narrow, over-optimized behaviors (clip-conv; Rl-plus; 8020entropy; deng2025trial; skywork; ReasonwithExplore; cheng2025reasoning). In response, recent work has focused on integrating entropy control mechanisms into the optimization process to preserve policy diversity. For instance, Cui et al. (clip-conv) regulate the impact of high-covariance tokens by applying a clipping function and a KL penalty. Meanwhile, Cheng et al. (cheng2025reasoning) augment the advantage function with a clipped, gradient-detached entropy term, which encourages deeper reasoning chains changing the original policy optimization direction.
@@ 2.2 Reinforcement Learning for LLM Agents
To enhance their autonomy, LLM agents are designed to interact with external environments using diverse toolsets (oaicua; tongyidr). However, training agents to complete multi-step tasks with these tools presents significant challenges for standard reinforcement learning, including sparse rewards and credit assignment problems. To address these issues, seminal works have introduced advanced training paradigms such as hierarchical RL (archer), autonomous learning (digirl), and off-policy Q-learning (digiq). Meanwhile, another line of research employs supervised fine-tuning (SFT) to directly enhance the models’ decision-making abilities, training them on vast datasets of high-quality tool-use trajectories to master complex environments and APIs (agentgym; zhang2024codeagent; qin2023toolllm). Recently, to leverage the training stability of GRPO (grpo), researchers have extended single-turn GRPO to multi-turn training settings with various training techniques to improve performance (jin2025search; gigpo; ragen). To guide learning in long-horizon scenarios, RLVMR (zhang2025rlvmr) introduces verifiable meta-reasoning rewards that provide dense, intermediate feedback on the agent’s reasoning process. However, these methods overlook the entropy dynamics unique to multi-turn settings. Because all turns share policy parameters, entropy adjustments cannot independently control early-turn exploration and late-turn exploitation. This coupling, combined with sparse rewards that provide no intermediate correction, causes the exploration-exploitation cascade failure, where severe entropy oscillations destabilize training and prevent effective long-horizon learning.
@@ 3.1 On-policy Optimization
where $A^{\pi_{\theta}}(s_{t},a_{t})=Q^{\pi_{\theta}}(s_{t},a_{t})-V^{\pi_{\theta}}(s_{t})$ is the advantage function. While the policy gradient provides an unbiased estimator of $\nabla_{\theta}J(\theta)$, directly optimizing this objective can lead to instability due to large policy updates. To address this, modern on-policy methods employ surrogate objective functions that approximate the policy gradient while ensuring stable learning. The standard policy gradient can be reformulated as the surrogate objective $L^{PG}(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta_{\text{old}}}}\left[\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}\hat{A}_{t}\right]$. However, this surrogate objective is only valid for infinitesimally small updates. Proximal Policy Optimization (PPO) (ppo) addresses this limitation by constraining the policy ratio through a clipped surrogate objective: $L^{CLIP}(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta_{\text{old}}}}\left[\min\left(r_{t}(\theta)\hat{A}_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}\right)\right]$ where $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}$ is the importance sampling ratio, $\mathbb{E}_{\tau\sim\pi_{\theta_{\text{old}}}}[\cdot]$ denotes expectation over trajectories sampled under $\pi_{\theta_{\text{old}}}$, and $\epsilon$ defines the trust region bounds. Group Relative Policy Optimization (GRPO) (grpo) extends PPO by modifying the advantage function computation. Instead of using standard advantages $\hat{A}_{t}$, GRPO employs group-relative advantages, computed as $\tilde{A}_{t}=\frac{R_{t}-\mu_{g}}{\sigma_{g}+\delta}$. where $R_{t}$ is the return from timestep $t$, $\mu_{g}$ and $\sigma_{g}$ are the mean and standard deviation of returns within group $g$, and $\delta$ is a small constant for numerical stability. This group-based normalization provides more stable gradient estimates.
@@ 3.2 Problem formulation
We formalize the multi-turn task as a sequential decision-making reinforcement learning problem. A single LLM agent $\pi_{\theta}$ executes a task through $T$ turns over a trajectory $\tau=(s_{0},a_{0},r_{0},\ldots,s_{T},a_{T},r_{T})$. The reward is sparse, with $r_{t}=0$ for all intermediate turns and only the final turn receiving the task outcome reward, such that the total return is $R(\tau)=\sum_{t=0}^{T}\gamma^{t}r_{t}=r_{T}$, where $r_{T}\in\{0,1\}$ represents the binary task outcome. In our experimental settings, specifically ALFWorld (alfworld) and SciWorld (sciworld), we assume an undiscounted formulation where $\gamma=1$. All turns within the same task share the final outcome reward, creating a credit assignment challenge across sequential turns. The multi-turn policy optimization differs from standard RL in that losses accumulate across all $T$ turns before parameter updates: $L^{MT}(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta_{\text{old}}}}\left[\mathbb{E}_{t\sim T}\left[\min\left(r_{t}(\theta)A_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)A_{t}\right)\right]\right]$ where $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}$ and $A_{t}$ represents the advantage estimate at turn $t$ within the multi-turn trajectory. $\mathbb{E}_{t}[\cdot]$ denotes expectation over timesteps within turn $t$, while the outer expectation is over trajectories sampled under $\pi_{\theta_{\text{old}}}$.